OpenStudy (anonymous):

Which of the following represents 2x - 5y + 15 = 0 written in slope-intercept form?

OpenStudy (texaschic101):

slope intercept form is " y = mx + b ". 2x - 5y + 15 = 0 we have to get -5y on one side, so we will add 5y to both sides 2x + 15 = 5y now we have to divide by 5 to get y by itself (2x + 15)/5 = y 2/5x + 3 = y which equals y = 2/5x + 3